Define derivatives for the following functions

a.) y= √3x^2+1/3t

b.) y= e^-x sin^2x

c.) y= sin2x / sin^2x
1
Expert's answer
2019-12-12T10:19:17-0500

1)  y= √3x2+1/3t  

Thus,

dy/dx=d/dx(√3x2+1/3t)=2√3x+0=2√3x,

Assume t is constant here.

2)y= e-x sin2x

dy/dx= e-x d/dx( sin 2x)+sin2x d/dx(e-x )

dy/dx=e-x d/dx (1-cos2x)/2+sin2xd/dx(e-x )

dy/dx=e-x sin2x-sin2x e-x


3) y= sin2x / sin2x

y=2sin2x/(1-cos2x)

dy/dx={(1-cos2x) d/dx(2sin2x)-2sin2x d/dx(1-cos2x)}/(1-cos2x)2

dy/dx={4(1-cos2x)cos2x-4sin2x sin2x}/(1-cos2x)2

dy/dx={4cos2x-4cos22x-4sin22x}/(1-cos2x)2

dy/dx=(4cos2x-4)/(1-cos2x)2

dy/dx=-4/(1-cos2x)